Hello, Can someone help me with this problem,I did get an answer but I not sure if it is correct. Thank you.
A card is picked from a standard deck of cards. Determine the odds against and the odds in favor of selectin…
a card greater than 5 (ace is low).
ANSWER:
There are 36 cards greater then 5
= 36/52
= This means that 36/52 reduces to 9/13
What are the odds in favor 36:52 or 9:13
What are the odds against 52:36 or 13:9
Thank you again.

Answer by Fombitz(13828)

Two things,
No, you need to put odds in terms of favorable outcomes:unfavorable outcomes.
Also, if Ace is low (1), there are 8 cards greater than 5 (6,7,8,9,10,J,Q,K) so then
Cards greater than 5 are 32.
Cards less than 5 are 20.
Odds in favor of selecting a card greater than 5 are 32:20 or 8:5.
Odds against selecting a card greater than 5 are 20:32 or 5:8.
Typically odds are presented in lowest terms.
.
.
.
Probabilities are given as favorable outcomes over total outcomes.
P(card>5)=32/52=8/13
P(card<=5)=5/13